OpenStudy (anonymous):

Frank has a batting average of 0.105. Write this decimal as a fraction in simplest form. a. 21/200 b. 1/10 c. 19/200 d. 11/100

OpenStudy (anonymous):

so the answer is?

OpenStudy (turingtest):

105/1000=21/200

OpenStudy (anonymous):

105 21 0.105 = ------ = ------- 1000 200

OpenStudy (anonymous):

the anwser is

OpenStudy (anonymous):

21/200

OpenStudy (turingtest):

if you can't figure that out now you are beyond help

OpenStudy (anonymous):

i.e a.)

OpenStudy (anonymous):

can you expalin the problem to me

OpenStudy (anonymous):

first write the decimal number as a fraction then reduce it to lowest terms by cancelling out common factors

OpenStudy (anonymous):

i know the anwser i just dont understand it i am not retarded

OpenStudy (anonymous):

thanks

OpenStudy (turingtest):

well decimals can always be written as whatever the number after the decimal is divided by 10^n ie.) .035=35/1000 .2=2/10 etc. here we have 0.105=105/1000 now you can reduce it like a regular fraction divide top and bottom by 5: 21/200